# Babel-ImageNet: Massively Multilingual Evaluation of Vision-and-Language Representations

### v1

@@ Abstract
Vision-and-language (VL) models with separate encoders for each modality (e.g., CLIP) have become the go-to models for zero-shot image classification and image-text retrieval. The bulk of the evaluation of these models is, however, performed with English text only: the costly creation of language-specific image-caption datasets has limited multilingual VL benchmarks to a handful of high-resource languages. In this work, we introduce Babel-ImageNet, a massively multilingual benchmark that offers (partial) translations of 1000 ImageNet labels to 92 languages, built without resorting to machine translation (MT) or requiring manual annotation. We instead automatically obtain reliable translations of ImageNext concepts by linking them – via shared WordNet synsets – to BabelNet, a massively multilingual lexico-semantic network. We evaluate 8 different publicly available multilingual CLIP models on zero-shot image classification (ZS-IC) for each of the 92 Babel-ImageNet languages, demonstrating a significant gap between English ImageNet performance and that of high-resource languages (e.g., German or Chinese), and an even bigger gap for low-resource languages (e.g., Sinhala or Lao). Crucially, we show that the models’ ZS-IC performance on Babel-ImageNet highly correlates with their performance in image-text retrieval, validating that Babel-ImageNet is suitable for estimating the quality of the multilingual VL representation spaces for the vast majority of languages that lack gold image-text data. Finally, we show that the performance of multilingual CLIP for low-resource languages can be drastically improved via cheap, parameter-efficient language-specific training. We make our code and data publicly available: https://github.com/gregor-ge/Babel-ImageNet
@@ 1 Introduction
CLIP models (Radford et al., 2021; Jia et al., 2021; Pham et al., 2021) have arguably become the most widely used vision-and-language (VL) models, owing popularity to efficient inference based on separate yet semantically aligned encoders for the two modalities. Their bi-encoder architecture makes them ideal for efficient image-text retrieval (Lin et al., 2014; Plummer et al., 2015) and zero-shot image classification (Radford et al., 2021). They can also produce input vectors for supervised tasks such as image generation (Rombach et al., 2022) or cross-modal reasoning (Eichenberg et al., 2022; Li et al., 2023).
@@
Motivated by the observation that performance on ImageNet classification translates well to performance in many other image tasks (Recht et al., 2019; Fang et al., 2023), CLIP models are typically evaluated on zero-shot image classification (ZS-IC), i.e., by comparing the representation of an image with text representations of class labels, whereby ImageNet (Deng et al., 2009) is the most prominent benchmark. With ImageNet class labels available only in English, this supports only evaluation of monolingual English models (i.e., models trained with English captions only). Although most CLIP models are trained on English-only image-caption data, some effort has been put into creating multilingual and monolingual non-English models by (1) training them from scratch (Bianchi et al., 2021; Ilharco et al., 2021; Yang et al., 2022; Jain et al., 2021) or (2) distilling them from English models (Carlsson et al., 2022; Chen et al., 2022; Zhang et al., 2022), typically using parallel data as supervision. Despite attempts to translate ImageNet labels to other languages (Bianchi et al., 2021; Yang et al., 2022), the language coverage remains very limited. Because of this, multilingual CLIP models have mainly been benchmarked on image-text retrieval datasets (Aggarwal and Kale, 2020; Bugliarello et al., 2022, inter alia), which predominantly cover only limited sets of mid-to-high resource languages.
Figure 1: Babel-ImageNet translations in 92 languages for the ImageNet class 309 (synset n02206856: “bee”). Font size of the language is proportional to the number of successfully translated ImageNet classes. [Image by Charles J. Sharp, CC BY-SA 3.0, via Wikimedia Commons]
@@
Rationale. Creating massively multilingual gold-standard datasets for VL tasks (e.g., image-text retrieval) is prohibitively expensive. Existing efforts (Aggarwal and Kale, 2020; Bugliarello et al., 2022; Thapliyal et al., 2022) either hire native speakers to write image captions in target languages or they resort to machine translation (MT) of English data, followed by manual post-editing by native speakers. The MT approach (the cheaper of the two), is, we argue, much more expensive for low-resource languages because (1) MT models are less accurate when translating to those languages, which implies a bigger post-editing effort for (2) bilingual annotators, native in the low-resource language and fluent in English: such annotators are more difficult to find for low-resource than for high-resource target languages (compare, e.g., Sinhala and German). In this work, we thus seek to create a robust massively multilingual benchmark for evaluating the quality of representation spaces of multilingual VL models, without resorting to MT or requiring any manual annotation effort. To be useful, such a benchmark needs to satisfy a crucial requirement: models’ performance across languages must be indicative of their performance for the same languages in tasks such as image-text retrieval, for which creating massively multilingual (gold-standard) evaluation datasets is too expensive.
@@
Contributions. With this in mind, we create Babel-ImageNet, a massively multilingual dataset for zero-shot image classification that offers (partial) translations of the 1000 ImageNet-1k classes to 92 languages. To obtain robust translations of ImageNet labels in other languages, we leverage the connection between ImageNet classes, which are derived from WordNet (Miller, 1994) synsets, and BabelNet (Navigli and Ponzetto, 2010), a massively multilingual lexico-semantic network, also (in part) derived from WordNet. Relying on the multilingual BabelNet synsets (and WordNet synset identifiers of ImageNet classes) to pivot between languages we avoid problems known to occur with machine translation of short phrases without context, e.g., due to polysemy [^1] For example, the ImageNet class walking stick refers to the insect and not the inanimate object.. Exploiting BabelNet allows us to automatically obtain labels for ImageNet concepts in many languages, removing the need for MT and manual annotation.
@@
We evaluate 8 different multilingual CLIP models on Babel-ImageNet, observing that all of them exhibit poor performance for low-resource languages. Crucially, we validate that Babel-ImageNet is a meaningful benchmark for measuring the quality of multilingual VL representations by comparing models’ performance on Babel-ImageNet against their performance on established multilingual image-text retrieval datasets. Finally, we propose a computationally efficient approach for improving multilingual CLIP models for low-resource languages. This modular language specialization approach yields large performance gains (>20% for some of the low-resource languages).
@@ 2 Related Work
We first provide an overview of existing benchmarks for evaluating multilingual VL models, followed by a brief overview of multilingual CLIP models, commonly used for efficient image-text retrieval.
2.1 Multilingual Vision-and-Language Benchmarks
@@
Early multilingual VL models (Gella et al., 2017; Wehrmann et al., 2019; Kim et al., 2020; Burns et al., 2020; Ni et al., 2021; Geigle et al., 2022; Zhou et al., 2021) were often evaluated in image-text retrieval on Multi30k (Elliott et al., 2016, 2017; Barrault et al., 2018), an extension of Flickr30k (Plummer et al., 2015) to German, French, and Czech, as well as on the Japanese (Yoshikawa et al., 2017) and Chinese (Li et al., 2019) translations of MSCOCO (Lin et al., 2014). More recent models were evaluated on multilingual image-text retrieval benchmarks: XTD (Aggarwal and Kale, 2020) (10 languages) and WIT (Srinivasan et al., 2021) (108 languages). Both these benchmarks, however, have prominent shortcomings. XTD predominantly contains examples from Karpathy’s training portion of MSCOCO (Karpathy and Fei-Fei, 2017), which is commonly used for pretraining of VL models: this constitutes a case of data leakage because XTD’s multilingual captions are directly translated from the original English captions (Bugliarello et al., 2022). WIT collects image-caption pairs from Wikipedia(s), which leads to two problems: (1) Wikipedia captions are abundant with named entity mentions, surface forms of which are often identical across a number languages – this artificially equates the difficulty of retrieval across languages (Zhai et al., 2022); (2) The number of image-caption pairs for a language depends on the size of its Wikipedia: this not only prevents direct performance comparisons across languages but also results in overly optimistic performance estimates for low-resource languages with merely a few hundred image-caption test pairs. More recently, IGLUE (Bugliarello et al., 2022) was introduced as the first benchmark to also include reasoning tasks like visual QA (Pfeiffer et al., 2022), primarily meant to test cross-encoder models that jointly encode image-text pairs (Ni et al., 2021; Zhou et al., 2021; Zeng et al., 2022). IGLUE also introduces the image-caption dataset xFlickrCo, a combination of Flickr30k and MSCOCO with captions in 7 languages. A recent dataset, XM3600 (Thapliyal et al., 2022), encompasses 3600 images (balanced by geography of origin) with captions in 36 languages. Despite high-quality [^2] The authors explicitly acknowledge the very high cost of human captioning of images in 36 languages (regretfully, they do not explicitly disclose the cost)., it has yet to become widely adopted – along with Zhai et al. (2023), we are among the first to use it.
@@
Motivated by monolingual CLIP models in other languages, translations of ImageNet classes have emerged for a handful of high-resource languages: Italian (obtained with MT) (Bianchi et al., 2021), Chinese (human translations) (Yang et al., 2022), and Japanese [^3] https://github.com/rinnakk/japanese-clip and Arabic [^4] https://github.com/LAION-AI/CLIP_benchmark/pull/68 (undisclosed translation methods). Extending ImageNet to more languages, however, is not feasible at scale: on the one hand, manual translation requires finding native speakers of low-resource languages (additionally fluent in English), which is challenging; machine translation of concepts (i.e., words and short phrases) out of context, on the other hand, is problematic due to polysemy – it thus also requires validation of the translated senses by bilingual annotators: this is especially critical for low-resource languages for which current MT systems are still lacking (Costa-jussà et al., 2022). In contrast, by exploiting the fact that ImageNet classes correspond to Wordnet synsets which, in turn, have corresponding massively multilingual synsets in BabelNet, we are able to create the first robust massively multilingual translation of ImageNet classes, avoiding the caveats of polysemy associated with automatic translation of concepts.
2.2 Multilingual CLIP
@@
While CLIP Radford et al. (2021) is not the first model for embedding images and text in the shared representation space, it has arguably become the most widely used one, owing its effectiveness – especially in ZS-IC – to the immense pretraining corpus. Older models, not exposed to large-scale VL pretraining, e.g., (Faghri et al., 2018) for English and (Gella et al., 2017; Wehrmann et al., 2019; Kim et al., 2020; Burns et al., 2020) multilingually, focused predominantly on text-to-image retrieval and were not shown to exhibit ZS-IC abilities. MURAL (Jain et al., 2021) was the first – albeit not publicly released – multilingual CLIP model, trained on billions of multilingual image-caption pairs. To the best of our knowledge, the only publicly available multilingual CLIP models trained “from scratch” are the OpenCLIP models (Ilharco et al., 2021) – pretrained using the full multilingual LAION5B dataset (Schuhmann et al., 2022), consisting of 5B image-caption pairs covering 100+ languages. Monolingual CLIP models for a few languages other than English (e.g., Italian, Chinese) have also been released (Bianchi et al., 2021; Yang et al., 2022): due to comparatively small pretraining data, they trail the English performance. Given the huge computational cost of training a multilingual CLIP from scratch, teacher distillation (Reimers and Gurevych, 2020) has become popular as an efficient alternative (Carlsson et al., 2022; Chen et al., 2022; Zhang et al., 2022): a pretrained multilingual text encoder (e.g., XLM-R Conneau et al. (2020)) is forced (commonly using parallel sentences) to align its representation space to the text encoder of English CLIP.
3 Babel-ImageNet: Massively Multilingual Zero-Shot Image Classification
@@
Why (massively) multilingual ZS-IC? With class labels in a particular language we can evaluate VL models in language-specific ZS-IC. Note that the goal is not to improve the image classification performance – using labels in any other language yields worse performance compared to using English labels. Instead, we argue that a model’s language-specific ZS-IC performance is a good estimate of the quality of its multilingual VL representation space for the language, and thus a good predictor of the model’s performance for that language in “real” tasks (e.g., image-caption retrieval).
@@
WordNet as a matchmaker for ImageNet and BabelNet. Unlike in most image classification datasets (e.g., CIFAR10, Oxford Pets (Parkhi et al., 2012), Flowers102 (Nilsback and Zisserman, 2008)), where image classes are words, ImageNet (Deng et al., 2009) links images to concepts, represented with sets of synonyms (synsets) from English WordNet (Miller, 1994). BabelNet (Navigli and Ponzetto, 2010) is a massively multilingual lexico-semantic network, automatically created by merging and consolidating numerous lexico-semantic resources: from WordNets in dozens of languages (e.g., Hamp and Feldweg (1997); Pianta et al. (2002)) to (massively multilingual) Wikipedia and WikiData Vrandečić (2012). Crucially for our efforts, BabelNet is (1) also organized in (multilingual) synsets, containing synonyms across many languages and (2) each of its synsets has an explicit link to the corresponding (English) WordNet synset (if such exists). With WordNet as the seam between ImageNet and BabelNet, we are able to create a massively multilingual ZS-IC benchmark, without resorting to manual annotation or MT.
Language Selection. While one could obtain ImageNet class translations in all BabelNet languages, [^5] The latest BabelNet version (5.2) covers 520 languages: https://babelnet.org/statistics. We found 298 of them present in at least 10 synsets that correspond to classes of ImageNet-1K; we release labels for these 298 languages with at least 10 classes. We also release our code, so that anyone with access to BabelNet can create translations in additional languages. we limit our evaluation to 92 non-English languages (counting unique ISO codes) covered by the pretraining corpora of XLM-R (Conneau et al., 2020). The motivation for this decision is twofold: (1) 50K images in ImageNet and up to 80K label + prompt combinations per language – to be embedded with each model included in our comparative evaluation – make even the ZS-IC evaluation computationally intensive, limiting the total number of languages for which we could carry it out with our computational resources; (2) the majority of publicly available multilingual CLIP models have been derived precisely from XLM-R as the initial multilingual text encoder (see Table 1).
@@
Class Label Translation and Cleaning Process. For each ImageNet class, we first fetch all word senses in any of the 92 languages and English from the corresponding BabelNet synset (using the WordNet synset ID for matching). We next remove senses from other languages for which an identical English sense exists. We do this because having target language labels identical to English labels would allow multilingual VL models to rely on their high-quality English representations. With those being semantically (much) better than representations of other languages, this would lead to misleadingly optimistic estimates of models’ multilingual abilities. On average, this removes $84\pm 39$ classes from a language-specific benchmark (i.e., we lose classes for which all available BabelNet senses in a language also exist as English senses in the same synset). Next, we eliminate all senses that were added to BabelNet via machine translation, removing the potential negative effects of context-agnostic MT from our benchmark. This mostly affects high(er)-resource languages and removes on average $148\pm 184$ classes. Finally, we select for every remaining class the first sense in the respective language (according to the order of senses in BabelNet) as our final language-specific class label. While the above process is not error-free, we estimate [^6] Based on manual inspection of class labels for a handful of languages that the authors speak. we estimate that less than 1% of labels are incorrect. We believe this to be a very acceptable error rate, considering (i) that around 6% of ImageNet images are mislabeled (Northcutt et al., 2021) in the first place and (ii) that there are also erroneous mappings between ImageNet images and (English) WordNet synsets (Nielsen, 2018; Radford et al., 2021).
@@
Grouping Languages in Evaluation. On the one hand, comparing models’ performance over 92 languages (+English) is unwieldy; averaging performance across all languages, on the other hand, is too reductive and consequently not particularly informative. We thus opt for the middle ground: we group Babel-ImageNet languages in three buckets based on their number of classes (less than 333, 334 to 667, and 668 to 1000). We argue that the number of classes is a reasonable proxy for general “resourceness” of a language (see §C.1 for the full list of languages and corresponding numbers of classes) and accordingly designate the three groups as low-, mid-, and high-resource, encompassing 41, 35, and 16 languages, respectively. Additionally, given that ZS-IC becomes easier with fewer classes, averaging results across languages with more comparable numbers of classes (i.e., within each of our groups) makes more sense than averaging them across all languages. Nonetheless, due to differing sets of classes, we caution against direct performance comparisons of results between groups or across individual languages. Instead, for any particular language or language group Babel-ImageNet allows for a direct comparison of competing multilingual VL models.
4 Comparative Evaluation of CLIP Models
Models. We briefly describe the public CLIP variants we benchmark on Babel-ImageNet (overview in Table 1). All of them encode images with Vision Transformers (ViT) (Dosovitskiy et al., 2021) albeit of different sizes and with differently sized input patches (e.g., B-32 = Base Transformer with 32$\times$32-pixel patches).
OpenAI: The original CLIP model (Radford et al., 2021), trained contrastively on 400M English-only image-caption pairs. The text encoder is trained from scratch (i.e., not initialized with any pretrained weights).
OpenCLIP: The OpenCLIP project (Ilharco et al., 2021) aims to replicate the OpenAI models using the public Laion datasets (Schuhmann et al., 2021, 2022). Two multilingual models have been trained on the the multilingual LAION5B dataset: the B-32 model with the text encoder initialized with the weights of XLM-R (Base) and the H-14 model with XLM-R-Large as the initialization of the text encoder. The B-32 variant is trained with the original contrastive CLIP objective, whereas the H-14 model was trained via locked image tuning (LiT, (Zhai et al., 2022)) in which the pretrained image encoder of the English H-14 OpenCLIP model is frozen and only the parameters of the text encoder are updated.
SentenceTransformer (ST): One of the first distillation-based multilingual CLIP-like models. It was obtained via the established distillation approach of Reimers and Gurevych (2020), [^7] https://huggingface.co/sentence-transformers/clip-ViT-B-32-multilingual-v1, using over 50M EN-X parallel sentences (X being one of 49 other languages) as supervision. They distill an (distilled, sentence-based) mBERT student (Devlin et al., 2019) from the English OpenAI B-32 teacher.
M-CLIP: Multilingual-CLIP (M-CLIP) Carlsson et al. (2022) is another model distilled using mBERT and OpenAI B-32. Unlike ST, they use automatic translations of 3M English image captions (from public image-text datasets) to 69 languages as parallel supervision for the distillation training. Post-publication, the authors additionally released a set of models [^8] https://github.com/FreddeFrallan/Multilingual-CLIP/blob/main/larger_mclip.md with XLM-R-Large (instead of mBERT) as the student initialization and (English) OpenAI B-32, L-14, and OpenCLIP B-16+ as teachers, respectively. These models were trained on 7M captions provided by Li et al. (2022), machine-translated to 48 languages. The original English captions were not used in distillation training.
AltCLIP: This multilingual CLIP model by Chen et al. (2022) distills an XLM-R-Large student with OpenAI L-14 as teacher, targeting 9 languages and using as training data a mix of machine-translated captions, multilingual captions sampled from LAION5B, and aligned English-X sentence pairs. After distillation, the authors fine-tune the model via LiT using sampled image-text pairs from LAION5B.
Source Text Image Objective #Data #Langs
OpenAI Base B-32 CLIP 400M 1
OpenCLIP XLM-R-Base B-32 CLIP 5B >100
OpenCLIP XLM-R-Large H-14 LiT 5B >100
M-CLIP mBERT B-32 distill 3M 69
M-CLIP XLM-R-Large B-32 distill 7M 48
M-CLIP XLM-R-Large B-16+ distill 7M 48
M-CLIP XLM-R-Large L-14 distill 7M 48
SentenceTransformer (distilled) mBERT B-32 distill >50M 49
AltCLIP XLM-R-Large L-14 distill+LiT 50M+100M 9

### v2

@@ Abstract
Vision-and-language (VL) models with separate encoders for each modality (e.g., CLIP) have become the go-to models for zero-shot image classification and image-text retrieval. They are, however, mostly evaluated in English as multilingual benchmarks are limited in availability. We introduce Babel-ImageNet, a massively multilingual benchmark that offers (partial) translations of ImageNet labels to 100 languages, built without machine translation or manual annotation. We instead automatically obtain reliable translations by linking them – via shared WordNet synsets – to BabelNet, a massively multilingual lexico-semantic network. We evaluate 11 public multilingual CLIP models on zero-shot image classification (ZS-IC) on our benchmark, demonstrating a significant gap between English ImageNet performance and that of high-resource languages (e.g., German or Chinese), and an even bigger gap for low-resource languages (e.g., Sinhala or Lao). Crucially, we show that the models’ ZS-IC performance highly correlates with their performance in image-text retrieval, validating the use of Babel-ImageNet to evaluate multilingual models for the vast majority of languages without gold image-text data. Finally, we show that the performance of multilingual CLIP can be drastically improved for low-resource languages with parameter-efficient language-specific training. We make our code and data publicly available: https://github.com/gregor-ge/Babel-ImageNet
@@ 1 Introduction
CLIP models (Radford et al., 2021; Jia et al., 2021; Pham et al., 2021) have become widely used vision-and-language (VL) models, owing popularity to efficient inference based on separate yet semantically aligned encoders for the two modalities. Their bi-encoder architecture makes them ideal for efficient image-text retrieval (Lin et al., 2014; Plummer et al., 2015) and zero-shot image classification (Radford et al., 2021), or to produce downstream features for supervised tasks such as image generation (Rombach et al., 2022) or cross-modal reasoning (Eichenberg et al., 2022; Li et al., 2023).
@@
Motivated by the observation that performance on ImageNet classification translates well to performance in many other image tasks (Recht et al., 2019; Fang et al., 2023), CLIP models are typically evaluated on zero-shot image classification (ZS-IC), i.e., by comparing the representation of an image with text representations of class labels, whereby ImageNet (Deng et al., 2009) is the most prominent benchmark. With ImageNet class labels available only in English, this supports only evaluation of monolingual English models (i.e., models trained with English captions only). Although most CLIP models are trained on English-only image-caption data, some effort has been put into creating multilingual and monolingual non-English models by (1) training them from scratch (Bianchi et al., 2021; Ilharco et al., 2021; Yang et al., 2022; Jain et al., 2021; Zhai et al., 2023) or (2) distilling them from English models (Carlsson et al., 2022; Chen et al., 2022b; Zhang et al., 2022; Visheratin, 2023), typically using parallel data as supervision. Despite attempts to translate ImageNet labels to other languages (Bianchi et al., 2021; Yang et al., 2022), the language coverage remains very limited. Because of this, multilingual CLIP models have mainly been benchmarked on image-text retrieval datasets (Aggarwal and Kale, 2020; Bugliarello et al., 2022; Thapliyal et al., 2022), which cover only limited sets of mid-to-high resource languages.
@@
Creating massively multilingual datasets for VL tasks (e.g., image-text retrieval) is prohibitively expensive. Existing efforts (Aggarwal and Kale, 2020; Bugliarello et al., 2022; Thapliyal et al., 2022) either hire native speakers to write image captions in target languages or resort to machine translation (MT) of English data, followed by manual post-editing by native speakers. The MT approach (the cheaper of the two), is, we argue, still too expensive for low-resource languages because MT models are less accurate when translating to those languages, which implies a bigger post-editing effort for bilingual annotators, native in the low-resource language and fluent in English; in addition, such annotators are more difficult to find for low-resource than for high-resource target languages. In this work, we thus seek to create a robust massively multilingual benchmark for evaluating the quality of representation spaces of multilingual VL models, without resorting to MT or requiring any manual annotation effort. To be useful, such a benchmark needs to satisfy a crucial requirement: models’ performance across languages must be indicative of their performance for the same languages in tasks such as image-text retrieval, for which creating massively multilingual (gold-standard) evaluation datasets is too expensive.
@@
Contributions. With this in mind, we create Babel-ImageNet, a massively multilingual dataset for zero-shot image classification that offers (partial) translations of the 1000 ImageNet classes to 100 languages. To obtain robust translations of ImageNet labels in other languages, we leverage the connection between ImageNet classes, which are derived from WordNet (Miller, 1994) synsets, and BabelNet (Navigli and Ponzetto, 2010), a massively multilingual lexico-semantic network, also (in part) derived from WordNet. Relying on the multilingual BabelNet synsets (and WordNet synset identifiers of ImageNet classes) to pivot between languages, we avoid problems known to occur with machine translation of short phrases without context, e.g., due to polysemy [^1] For example, the ImageNet class walking stick refers to the insect and not the inanimate object.. Exploiting BabelNet allows us to automatically obtain labels for ImageNet concepts in many languages, removing the need for MT and manual annotation.
@@
We evaluate 11 different multilingual CLIP models on Babel-ImageNet, observing that all of them exhibit poor performance for low-resource languages. Crucially, we validate that Babel-ImageNet is a meaningful benchmark for measuring the quality of multilingual VL representations by comparing models’ performance on Babel-ImageNet against their performance on established multilingual image-text retrieval datasets. Babel-ImageNet thus allows us to evaluate models in languages not covered by those datasets and it additionally expands the retrieval-focused evaluation with the ZS-IC task in languages included in the established datasets. Finally, we propose a computationally efficient approach for improving multilingual CLIP models for low-resource languages. This modular language specialization approach yields large performance gains ($>$20% for some of the low-resource languages).
Figure 1: Illustrating the creation of Babel-ImageNet: ImageNet classes correspond to WordNet IDs, which are integrated into BabelNet, a multilingual semantic net. Through this, we look up synonymous word senses in all available languages, perform some cleaning and filtering and select one sense as label.
@@ 2 Related Work
Multilingual Vision-and-Language Benchmarks.
@@
Early multilingual VL models (Gella et al., 2017; Wehrmann et al., 2019; Kim et al., 2020; Burns et al., 2020; Ni et al., 2021; Geigle et al., 2022; Zhou et al., 2021) were often evaluated in image-text retrieval on Multi30k (Elliott et al., 2016, 2017; Barrault et al., 2018), an extension of Flickr30k (Plummer et al., 2015) to German, French, and Czech, as well as on the Japanese (Yoshikawa et al., 2017) and Chinese (Li et al., 2019) translations of MSCOCO (Lin et al., 2014). More recent models were evaluated on multilingual image-text retrieval benchmarks: XTD (Aggarwal and Kale, 2020) (10 languages) and WIT (Srinivasan et al., 2021) (108 languages). Both these benchmarks, however, have prominent shortcomings. XTD predominantly contains examples from Karpathy’s training portion of MSCOCO (Karpathy and Fei-Fei, 2017), which is commonly used for pretraining of VL models, which constitutes a case of data leakage (Bugliarello et al., 2022). WIT collects image-caption pairs from Wikipedia(s), which are abundant with named entity mentions that are often identical across a number languages – this artificially equates the difficulty of retrieval across languages (Zhai et al., 2022). More recently, IGLUE (Bugliarello et al., 2022) was introduced as the first benchmark to also include reasoning tasks like visual QA (Pfeiffer et al., 2022), primarily meant to test cross-encoder models that jointly encode image-text pairs (Ni et al., 2021; Zhou et al., 2021; Zeng et al., 2022). IGLUE also introduces the image-caption dataset xFlickrCo, a combination of Flickr30k and MSCOCO with new captions in 7 languages. Another recent dataset, XM3600 (Thapliyal et al., 2022), encompasses 3600 images (balanced by geography of origin) with captions in 36 languages written from scratch by native speakers.
@@
Motivated by monolingual CLIP models in other languages, translations of ImageNet classes have emerged for a handful of high-resource languages: Italian (obtained with MT) (Bianchi et al., 2021), Chinese (human translations) (Yang et al., 2022), and Japanese222github.com/rinnakk/japanese-clip and Arabic333github.com/LAION-AI/CLIP_benchmark/pull/68 (undisclosed methods). Extending ImageNet to more languages, however, is not feasible at scale, due to the challenges of finding native speakers to translate or verify machine-translated labels. With Babel-ImageNet, we exploit the linkage between ImageNet and BabelNet through WordNet, to create the first robust massively multilingual translation of ImageNet classes, avoiding the caveats of polysemy associated with automatic translation of concepts.
Multilingual CLIP.
@@
CLIP Radford et al. (2021) is not the first model to embed images and text in a shared representation space, but it has arguably become the most widely used one, owing its effectiveness – especially in ZS-IC – to the immense pretraining corpus. Older models, not exposed to large-scale VL pretraining, e.g., (Faghri et al., 2018) for English and (Gella et al., 2017; Wehrmann et al., 2019; Kim et al., 2020; Burns et al., 2020) multilingually, focused predominantly on text-to-image retrieval and were not shown to exhibit ZS-IC abilities. MURAL (Jain et al., 2021) was the first – albeit not publicly released – multilingual CLIP model, trained on billions of multilingual image-caption pairs. The only publicly available multilingual CLIP models trained “from scratch” are the OpenCLIP models (Ilharco et al., 2021) (trained on the full multilingual LAION5B dataset (Schuhmann et al., 2022)) and mSigLIP Zhai et al. (2023) (trained on Google’s WebLI Chen et al. (2022a)). Monolingual CLIP models for a few languages other than English (e.g., Italian, Chinese) have also been released (Bianchi et al., 2021; Yang et al., 2022): due to comparatively small pretraining data, they trail the English models. Given the huge computational cost of training a multilingual CLIP from scratch, teacher distillation (Reimers and Gurevych, 2020) has become popular as an efficient alternative (Carlsson et al., 2022; Chen et al., 2022b; Zhang et al., 2022): a multilingual text encoder (e.g., XLM-R Conneau et al. (2020)) is forced (commonly using parallel sentences) to align its representation space to the text encoder of English CLIP.
3 Babel-ImageNet
Why (massively) multilingual ZS-IC?
@@ 3 Babel-ImageNet
With class labels in a particular language we can evaluate VL models in language-specific ZS-IC. Note that the goal is not to improve the classification performance – using labels in any other language yields worse performance compared to using English labels. Instead, we argue that a model’s language-specific ZS-IC performance is a good estimate of the quality of its multilingual VL representation space for the language, and thus a good predictor of the model’s performance for that language.
In addition, prior work evaluated models mainly with retrieval. For the languages, where such data exists, we provide a more comprehensive evaluation: ImageNet covers a far more diverse set of concepts than image captions usually contain.
WordNet as a matchmaker for ImageNet and BabelNet.
@@
Unlike in most image classification datasets (e.g., CIFAR10, Oxford Pets (Parkhi et al., 2012), Flowers102 (Nilsback and Zisserman, 2008)), where image classes are words, ImageNet (Deng et al., 2009) links images to concepts, represented with sets of synonyms (synsets) from English WordNet (Miller, 1994). BabelNet (Navigli and Ponzetto, 2010) is a massively multilingual lexico-semantic network, automatically created by merging and consolidating numerous lexico-semantic resources: from WordNets in dozens of languages (e.g., Hamp and Feldweg (1997); Pianta et al. (2002)) to (massively multilingual) Wikipedia and WikiData Vrandečić (2012). [^4] BabelNet v5.2 consolidates 53 sources Crucially for our efforts, BabelNet is (1) also organized in (multilingual) synsets, containing synonyms across many languages and (2) each of its synsets has an explicit link to the corresponding (English) WordNet synset (if such exists). With WordNet as the seam between ImageNet and BabelNet, we are able to create a massively multilingual ZS-IC benchmark, without resorting to manual annotation or MT.
Class Translation and Cleaning Process.
@@
We illustrate the general process in Figure 1. For every ImageNet class, using its WordNet synset ID, we fetch all synonyms in every language including English (where available). Next, we remove words that match an English word (of the same class) because models use their high-quality English representations for these words, distorting results and leading to misleadingly optimistic estimates of models’ multilingual abilities. Next, we eliminate all words that were added to BabelNet via machine translation, removing the potential negative effects of context-agnostic MT from our benchmark. Finally, we select for every class in every language the first remaining word (according to the order in BabelNet) as our final language-specific class label. Classes with no word are removed.
Language Selection.
We find that of the 520 languages in BabelNet v5.2, 298 have at least 10 classes using our process, the majority of them being low- and very-low resource languages. We limit our evaluation to 100 non-English languages for more balanced selection of low-, mid- and high-resource languages. To this end, we combine the 92 languages (counting unique ISO codes) covered by the pretraining corpora of XLM-R (Conneau et al., 2020) with 8 manually selected languages not covered by machine translation (neither Google, Microsoft, nor by NLLB Costa-jussà et al. (2022)).
Grouping Languages in Evaluation.
@@
Comparing models’ performance over 100 languages (+English) is still unwieldy but averaging performance across all languages is too reductive and consequently not particularly informative. We thus opt for the middle ground: we group Babel-ImageNet languages in four buckets based on their number of classes ($<$101, 101 to 333, 334 to 667, and 668 to 1000). We argue that the number of classes is a reasonable proxy for general “resourceness” of a language (see §B.1 for the full list of languages and corresponding numbers of classes) and accordingly designate the four groups as very-low-, low-, mid-, and high-resource, encompassing 17, 32, 35, and 16 languages, respectively. Additionally, given that ZS-IC becomes easier with fewer classes, averaging results across languages with more comparable numbers of classes makes more sense than averaging them across all languages.
Verification.
BabelNet mappings are automatically created and not error-free even though, for example, over 90% of Wikipedia-WordNet mappings were manually validated with a precision over 99.5% Navigli et al. (2021). Still, we also manually verify our mappings on 4 languages with native speakers [^5] The authors and colleagues. Languages: de, ur, tr, hr and on the 3 smallest languages (xh, om, ha) with online dictionaries; we find on average 5.4% errors. We believe this to be a very acceptable error rate, considering (i) that around 6% of ImageNet images are mislabeled (Northcutt et al., 2021) and (ii) that there are also erroneous mappings between ImageNet images and WordNet synsets (Nielsen, 2018; Radford et al., 2021).
4 Benchmarking CLIP Models
Source Objective #Data #Langs
OpenAI CLIP 400M 1
OpenCLIP CLIP/LiT 5B $>$100
M-CLIP distill 3M 69
M-CLIP distill 7M 48
SentenceTransformer distill $>$50M 49
AltCLIP distill+LiT 50M+100M 9
SigLIP CLIP 900M 100
NLLB-SigLIP LiT 20M 201